this year jack's age is three times tom's age . in six year's time jack will only be 2 times older than tom what is jack's ages this year
Answer by ikleyn(52777)

.
this year jack's age is three times tom's age .
in six year's time jack will only be 2 times older than tom
what is jack's ages this year
~~~~~~~~~~~~~~~~~~~

x  = Tom's age this years

3x = Jack's age.


In 6 years time

    3x + 6 = 2*(x+6).


Simplify and find x

    3x + 6  = 2x + 12

    3x - 2x = 12 - 6

       x    =    6


ANSWER.  Jack is 3x = 18 years old this year.

Solved.